OpenStudy (anonymous):

This is hard, help, medal and fan. A waiter can find his/her gross pay for the week, with a 15% tip included, by using this formula: Gross Pay = (Hourly Wage)·(Hours Worked) + (0.15)·(Cost of Meals) Suppose that Serena worked 35 hours at $3.60 an hour and served $1210 in meals. How much did she make that week? Select one: a. $1213.60 b. $1517.50 c. $289.60 d. $307.50

OpenStudy (whpalmer4):

Gross pay = (35 hours)($3.60/hour) + (0.15)*($1210) Get out the calculator...

OpenStudy (kittyjosh):

It is simple, just do Gross Pay = (Hourly Wage)·(Hours Worked) + (0.15)·(Cost of Meals) Gross Pay = (3.60)*(35)+(0.15)*(1210) And you will find the answer to be D. $307.50

OpenStudy (anonymous):

i got 853, so obviously i did something wrong. haha thanks guys.

OpenStudy (whpalmer4):

You should figure out what you did wrong! Otherwise, you'll do it again, and might not realize it. 35*3.60 = 126 0.15*1210 = 181.50 126+181.50 = 307.50

OpenStudy (anonymous):

i accidentally multiplied 0.6 instead of 0.15. lol whoops

OpenStudy (whpalmer4):

Yep, that would do it!